# LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data

### v1

@@ 4.3. Application: Search & Ranking
4.3.1. Dataset
We evaluate LOTUS’ performance on the search and ranking task using three datasets, including BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, as well as two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generate to evaluate more complex, reasoning-based ranking criteria over the data. For each, we report nDCG@10, as well as the execution time (ET) in seconds.
The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance given each scientific claim. We sample 300 scientific facts for our evaluation, and report the average ranking execution time across these samples.
While the SciFact dataset provides a sorting task based on a simple relevance criterion, our newly proposed benchmarks provide a more complex sorting criteria over a corpus of paper abstracts. Specifically, the ranking task is to find the papers that report the highest accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench respectively.
To generate CIFAR-bench, we took 100 abstracts from the Papers with Code Dataset (pap, [n. d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain the top-10.
We then synthetically generated HellaSwag-bench by prompting Llama-70B to create 200 paper abstracts, each with a specified accuracy value, randomly sampled from . This setup allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth. We note that an alternative approach to these tasks could efficiently leverage sem_map to extract accuracy values on abstracts from either dataset, then perform a structured sort. However, our evaluation focuses on assessing the semantic ranking capabilities of LOTUS’ top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may likely generalize to a wider set of reasoning-based sorting queries, such as ”which paper makes the most outrageous claim”, for which ground truth is less objective to evaluate. For CIFAR-bench and HellaSwag-bench we report results for trials of the ranking task, at temperature , similar to prior works (Khattab et al., 2023)
| Method | nDCG@10 | Execution Time (s) |
| Semantic Search | 0.712 | 0.009 |
| Semantic Search + Reranker | 0.741 | 2.64 |
| LOTUS Top-k (quickselect + sem-index) - Llama-70B | 0.775 | 33.6 |
| LOTUS Top-k (quickselect + sem-index) - GPT-4o | 0.800 | 11.2 |
@@ 4.3.1. Dataset
| nDCG@10 | Execution Time (s) | nDCG@10 | Execution Time (s) | |
| Semantic Search | 0.252 | 0.008 | 0.119 | 0.008 |
| Semantic Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
| LOTUS Top-k (quickselect) - Llama 70-B | 0.746 | 41.3 | 0.909 | 59.1 |
4.3.2. Baselines
We consider two simple baselines. The first baseline performs semantic search, as shown in Figure 30, using the E5Model (Wang et al., 2024) for retrieval. The second baseline performs search with re-ranking, as the pseudocode shows in Figure 31, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking.
4.3.3. LOTUS Programs
The proposed LOTUS program performs a semantic top- over the documents, as shown in Figure 32, which shows example pseudocode for the CIFAR-bench dataset. The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets, as shown in the figure. We note that for the SciFact dataset, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk. We report results using both Llama-70B and GPT-4o as the LM.
4.3.4. Results
@@
Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- operator for complex search tasks. In addition, Table 6 and Figure 33 highlight the rich implementation design space that this task presents. We walk through several note-worthy findings.
@@
We first turn our attention to Table LABEL:tab:ranking_scifact, which presents the results for each bench-marked method on the SciFact dataset, which uses a relevance-based sorting criterion. As expected, both semantic search programs with and without re-ranking present strong baselines. Specifically, the re-ranker model, which is a supervised model trained specifically for the task of relevance-based ranking, increases nDCG@10 by 3 percentage points, while trading off query efficiency compared to the semantic search baseline. Notably, the unsupervised LM-based LOTUS programs outperform the supervised re-ranker’s result quality. The table shows LOTUS semantic top-k program with Llama-70B and GPT-4o, which outperform the semantic search baseline by 6 and 9 percentage points respectively. The LOTUS programs with Llama-70B and GPT-4o also outperform the re-ranker by 3 and 5 points respectively, improving upon the quality of the supervised baseline. As expected, the improved result quality comes at a trade-off to query efficiency due to the cost of LM calls. Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipeline and trade-off result quality and efficiency depending on application-specific requirements.
@@
Turning our attention to Table 5, we study LOTUS’ generality in supporting arbitrary language-based ranking criteria over the dataset. On the CIFAR-bench and HellaSwag-bench datasets, which use a complex sorting criteria, we see that both semantic search baselines with and without re-ranking provide poor result quality with consistently low nDCG@10. The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with 49.4 and 44.8 points higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively. These accuracy gains reflect the powerful reasoning capabilities of LMs efficiently orchestrated over the data. As expected, these significant accuracy gains come at an increase to execution time.
@@
We now analyze the efficiency of LOTUS semantic top- implementation along with it’s proposed optimizations. Table 6 compares several semantic top- algorithms, namely a quick-select top- algorithm, a quick-select top- that leverages the similarity index for pivot selection, a heap-based top- algorithm, and a quadratic sorting algorithm. First, we see that the quadratic algorithm, which performs an LM comparison between each pair of input documents, offers consistently high result quality across each dataset. However, this method is prohibitively expensive, requiring more LM calls and over higher execution time than the alternative implementations. The heap top- and quick-select top- methods offer comparable result quality, but with interesting trade-offs in query efficiency. Notably the quick-select top- method offers lower execution time than the heap-based sorting method across all datasets, despite requiring more LM calls in some cases. This is because the quick-select top- implementation allows for efficient batch-processing in each round of the algorithm, whereas the heap-based top- incurs sequential LM calls during heap updates. For this reason, our current implementation leverages the quick-select top- algorithm, although we envision future iterations may leverage multiple algorithms and allow the user to declaratively trade-off accuracy, query throughput, and cost.
@@
In addition to providing an efficient top- algorithm, the table also demonstrates the use of LOTUS’ similarity index for optimizing top- query performance. The quick-select top- algorithm optimized with the semantic similarity index for pivot selection demonstrates lower execution time at no accuracy loss on SciFact, where the ranking criteria correlates likely with semantic similarity. On the other hand, for the CIFAR-bench and HellaSwag-bench datasets, where the ranking criteria does not correlate with semantic similarity, we see that the similarity index has no significant impact on the accuracy or efficiency top- performance.
Lastly, we analyze the impact of LOTUS’ cascade optimization applied to the quickselect top- implementation. Figure 33 compares the nDCG and execution time of implementing the operator with a single model, either Llama 8B or Llama 70B, to the implementation that leverages these models together using model cascades. We vary the confidence threshold of the cascade to generate several points in the trade-off space. We find that the cascade optimization offers diverse performance trade-offs, which can outperform the single-model oracle baseline, which uses Llama-70B. For instance, one cascade along the Pareto-frontier improves accuracy of the Llama-70B baseline by , while reducing execution time by , demonstrating substantial opportunities for automatically optimizing LOTUS’ semantic query pipelines.
@@ 5. Related Work
Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks or applications.
Palimpzest (Liu et al., 2024b) presents a declarative approach to data cleaning and extract-transform-load (ETL) tasks. The authors propose to automatically optimize relational operators with LM calls, and implement two row-wise relational operators, a newly proposed convert operator, which can be transparently optimized to perform entity extraction using LLMs, and an AI-based filter operation, logically similar to LOTUS’ sem_filter. The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis to implement user queries, and proposes several others as future work.
@@
LOTUS, in contrast to these prior works, defines a general-purpose programming model designed to capture broad-ranging applications with a core set of composable semantic operators, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions. In LOTUS’ current implementation, users can use sem_map to perform entity extraction over unstructured text fields, although future work may provide native functionality for entity extraction by integrating systems or optimizations of prior work. Additionally, we believe that several optimizations leveraged in prior work, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
LLM UDFs
Recent research work (Liu et al., 2024a) and existing analytical database vendors, such as
Google BigQuery (ver, [n. d.]),
Databricks (dat, [n. d.]) and
@@
AWS Redshift (noa, 2023a), alternatively offer LLM user-defined functions (UDFs) to the programmer. The LLM UDF programming model provides a lower-level interface, which is limited to logically row-wise LLM execution over the data, equivalent to LOTUS’ sem_map. In contrast, LOTUS’ programming model is declarative and provides a rich set of semantic operators, allowing the system to automatically orchestrate the LM to serve a variety of complex query patterns, including aggregations, ranking, and joins.
@@
Liu et al. (Liu et al., 2024a) study how to optimize LLM UDF functions, demonstrating performance gains with a de-duplication method and a prefix-sharing maximization method that reorders rows and re-formulates parameterized prompts to maximize key-value (KV) cache reuse during query execution. We believe these methods could be effective optimizations in future work at LOTUS’ batched execution layer.
@@
LM Programming Frameworks Recent LM programming frameworks, such as LangChain (lan, [n. d.]), LlamaIndex (lla, [n. d.]), and DSPy (Khattab et al., 2023), have gained significant popularity. These systems provide a set of abstractions for programming with LMs, including utilities for handling prompts and post-processsing LM outputs, and support for common use cases, such as RAG, chat-bots, and function-calling. DSPy focuses on abstracting LM pipelines as programs and provides automatic prompt optimization. In contrast to these systems, LOTUS’ programming model is designed for tasks involving bulk processing data with LLMs. While some of these systems support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized.
@@
ML-based Query Processing Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique opportunities for system design and optimization. MADLib (Hellerstein et al., 2012) extends SQL with efficient abstractions for supervised learning, unsupervised learning, and descriptive statistics. Prior works such as NoScope (Kang et al., 2017), TASTI (Kang et al., [n. d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML models over large datasets, typically in video analytics. Some optimizations, such as model cascades and predicate re-ordering, which were useful in these works are likewise useful for optimizing LOTUS pipelines with language models. However, our setting with natural-language reasoning tasks has new operators with significantly different semantics, such as sem_topk and sem_agg, requiring a new programming model as well as new query execution algorithms and optimizations.
@@ 6. Conclusion
In this work, we proposed semantic operators to provide the first declarative and general-purpose interface to serve bulk-semantic processing. We implement these operators in the LOTUS system to seamlessly extend the relational model and allow programmers to easily compose powerful reasoning-based query pipelines over vast corpora of structured and unstructured data. Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of LOTUS’ programming model as well as the efficiency and optimization opportunities of LOTUS’ query engine. For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized to achieve higher performance than existing implementations.

### v2

@@ 5.3. Application: Search & Ranking
Dataset. We evaluate LOTUS’ performance on the search and ranking task using three datasets: BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, and two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generated to evaluate more complex ranking criteria over the data. For each, we report nDCG@10 and execution time (ET) in seconds.
The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance to a given claim. We sample 300 claims for our evaluation, and report the average ranking execution time across these samples.
While the SciFact dataset provides a sorting task based on a simple relevance criterion, our generated benchmarks provide a more complex sorting criteria each over a corpus of paper abstracts. The task is to rank papers according to highest reported accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench, respectively.
To generate CIFAR-bench, we took 100 abstracts from the Papers with Code Dataset (pap, [n. d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain ground truth.
We then synthetically generated HellaSwag-bench by prompting Llama-70B to create 200 paper abstracts, each with a specified accuracy value, randomly sampled from . These datasets allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth an ranking criteria other than relevance-based ones. We note that an alternative approach to these tasks could efficiently leverage sem_map to extract accuracy values on abstracts from either dataset, then perform a structured sort. However, our evaluation focuses on assessing the semantic ranking capabilities of various top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may generalize to a wider set of reasoning-based ranking queries. We report results for trials at temperature , similar to prior works (Khattab et al., 2023)
| Method | nDCG@10 | ET (s) |
| Search | 0.712 | 0.009 |
| Search + Reranker | 0.741 | 2.64 |
| LOTUS Top-k - Llama-70B | 0.775 | 33.6 |
| LOTUS Top-k - GPT-4o | 0.800 | 11.2 |
@@
| nDCG@10 | ET (s) | nDCG@10 | ET (s) | |
| Search | 0.252 | 0.008 | 0.119 | 0.008 |
| Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
| LOTUS Top-k - Llama 70B | 0.710 | 39.6 | 0.975 | 63.6 |
Baselines. We consider two simple baselines: a semantic search, using the E5Model (Wang et al., 2024) for retrieval, and a search with re-ranking, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking. We use choose these models because they are recent high quality embedding and re-rankers.
LOTUS Program. Our LOTUS program performs a semantic top- over the documents.
The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets. For SciFact, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk. We report results using both Llama-70B and GPT-4o as the LM.
@@
Results. Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- for complex search tasks. In addition, Table 6 highlights the rich implementation design space that this task presents. We walk through several note-worthy findings. We first turn our attention to Table LABEL:tab:ranking_scifact, which presents the results for each bench-marked method on the SciFact dataset, which uses a relevance-based sorting criterion. As expected, both semantic search programs with and without re-ranking present strong baselines. Specifically, the re-ranker model, which is a supervised model trained specifically for the task of relevance-based ranking, increases nDCG@10 by 3 percentage points, while trading off query efficiency compared to the semantic search baseline. Notably, the unsupervised LM-based LOTUS programs outperform the supervised re-ranker’s result quality. The table shows LOTUS semantic top-k program with Llama-70B and GPT-4o, which outperform the semantic search baseline by 6 and 9 percentage points respectively. The LOTUS programs with Llama-70B and GPT-4o also outperform the re-ranker by 3 and 5 points respectively, improving upon the quality of the supervised baseline. As expected, the improved result quality comes at a trade-off to query efficiency due to the cost of LM calls. Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipelines and trade-off result quality and efficiency depending on application-specific requirements.
@@
Turning our attention to Table 5, we study LOTUS’ generality in supporting arbitrary language-based ranking criteria over the dataset. On the CIFAR-bench and HellaSwag-bench datasets, which use a complex sorting criteria, we see that both semantic search baselines with and without re-ranking provide poor result quality with consistently low nDCG@10. The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with and higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively. These accuracy gains reflect the powerful reasoning capabilities of LMs efficiently orchestrated over the data. As expected, these significant accuracy gains come at an increase to execution time.
@@
We now analyze the efficiency of our semantic top- implementation along with it’s proposed optimizations compared to alternatives. Table 6 compares several semantic top- algorithms, namely a quick-select top- algorithm, a quick-select top- that leverages the similarity index for pivot selection, a heap-based top- algorithm, and a quadratic sorting algorithm. First, we see that the quadratic algorithm, which performs an LM comparison between each pair of input documents, offers consistently high result quality across each dataset. However, this method is prohibitively expensive, requiring more LM calls and over higher execution time than the alternative implementations. The heap top- and quick-select top- methods offer comparable result quality, but with interesting trade-offs in query efficiency. Notably the quick-select top- method offers lower execution time than the heap-based sorting method across all datasets, despite requiring more LM calls in some cases. This is because the quick-select top- implementation allows for efficient batch-processing in each round of the algorithm, whereas the heap-based top- incurs sequential LM calls during heap updates. For this reason, our current implementation leverages the quick-select top- algorithm, although we envision future iterations may leverage multiple algorithms and allow the user to declaratively trade-off accuracy, query throughput, and cost.
@@
In addition to providing an efficient top- algorithm, the table also demonstrates our use of the similarity index for optimizing top- query performance. The quick-select top- algorithm optimized with the semantic similarity index for pivot selection demonstrates lower execution time at no accuracy loss on SciFact, where the ranking criteria correlates likely with semantic similarity. On the other hand, for the CIFAR-bench and HellaSwag-bench datasets, where the ranking criteria does not correlate with semantic similarity, we see that the similarity index has no significant impact on the accuracy or efficiency top- performance.
5.4. Application: ArXiv Paper Analysis
Dataset.
Lastly, we evaluate LOTUS’ use in analyzing a large corpus of documents, specifically focusing on the performance of sem_group_by and its approximation algorithm with accuracy guarantees. The dataset consists of 647 recent ArXiv articles we scraped, from the database (cs.DB), information retrieval (cs.IR), cryptography and security (cs.CR), and robotics (cs.RO) domains. The task is to group the papers according to the topics they discuss.
Baselines. The sem_group_by consists of two sub-tasks: (a) discovering the groups and labels, and (b) classifying each document using the discovered labels. As a baseline for the latter, we consider a semantic similarity join between the labels and each abstract. We use the E5 embedding model for retrieval.
IRIS Programs.
We use the LOTUS sem_group_by operator to find 5 groups based on the topic abstracts. We use Llama-70B as the LM, the E5Model as the retrieval model, and 100 samples for the approximation algorithm described in Section 4.
| () Advancements in Recommender Systems and Multimodal Data Integration |
| () Advancements in Generative Information Retrieval Systems |
| () Advancements in Large Language Models for Various Applications |
| () Advancements in AI Security and Malware Detection Techniques |
| () Advancements in Robotic Navigation and Manipulation Techniques |
Results. We first qualitatively analyze the labels discovered, in the first phase of LOTUS’ implementation. As Figure LABEL:tab:groupby shows, the discovered labels intuitively align with the ArXiv domains included in the paper scrap, including important and timely topics related to recommendation systems, retrieval, LLM applications, AI security, and robotics. Performing this unsupervised group discovery took 44.03 seconds, representing a tractable and efficient implementation.
Finally, turning our analysis to the classification subprocedure in our sem_group_by implementation, we evaluate the declarative algorithmic approximation. Figure 11 compares the execution time and classification accuracy with respect to the oracle implementation for the operator implemented with varying accuracy targets using . We see that the oracle implementation, at , offers the slowest execution time. On the other hand, the similarity-join baseline has faster execution time, but only accuracy with respect to the LM. We find our proxy-based optimization offers diverse performance trade-offs and allows the user to declaratively interpolate between these two extremes. We also note that the sampling procedure used to perform this optimization took less than 5 seconds, representing a small relatives cost.
@@ 6. Related Work
Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks.
Palimpzest (Liu et al., 2024b) presents a declarative approach to data cleaning and extract-transform-load (ETL) tasks. The authors propose to automatically optimize relational operators with LM calls, and implement two row-wise relational operators, a newly proposed convert operator, which can be transparently optimized to perform entity extraction using LLMs, and an AI-based filter operation, logically similar to LOTUS’ sem_filter. The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis and proposes several others as future work.
@@
In contrast to these works, semantic operators are a general-purpose query model designed to capture broad-ranging applications, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions, as an extension of the relational model. For example, users can use sem_map for row-wise entity extraction over unstructured text fields, but can also perform other operations over entire tables, such as sem_join.
We also propose multiple execution algorithms and novel optimizations for these bulk operators (Section 4).
Some of the optimizations in prior specialized systems, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
LLM UDFs.
Recent research (Liu et al., 2024a) and some existing DBMS vendors, such as
BigQuery (ver, [n. d.]),
@@
Databricks (dat, [n. d.]) and Redshift (noa, 2023), offer LLM user-defined functions (UDFs) in SQL. LLM UDFs are a lower-level programming interface that is limited to row-wise LLM execution over the data, equivalent to our sem_map semantic operator. In contrast, our semantic operator programming model is high-level and declarative, allowing the system to automatically select which an algorithm to use and manage calls to the LM to support complex operations, including aggregations, ranking, and joins.
@@
Liu et al. (Liu et al., 2024a) study how to optimize LLM UDF calls, demonstrating performance gains with a de-duplication and prefix-sharing method that reorders rows and input fields to maximize key-value (KV) cache reuse during LLM calls. This is orthogonal to our optimizations but could likely accelerate batch LLM calls from LOTUS.
@@
LM Programming Frameworks. LM programming frameworks such as LangChain (lan, [n. d.]), LlamaIndex (lla, [n. d.]), and DSPy (Khattab et al., 2023), have gained significant popularity. They provide abstractions for programming with LMs, including utilities for creating prompts and post-processing LM outputs, and modules for common calling patterns such as RAG, chat-bots, and agents. DSPy also optimizes a pipeline of LLM calls to maximize a target metric by tuning prompts and model weights. However, all these frameworks focus on expressing a pipeline of LLM calls that takes a single small input (or a stream of inputs for chat), as opposed to programs for processing large datasets. While some of them support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized. In contrast, semantic operators and LOTUS are designed for bulk semantic processing and introduce novel operators and optimizations for this task.
@@
ML-based Query Processing. Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique usage patterns and optimization opportunities. MADLib (Hellerstein et al., 2012) extends SQL with abstractions for machine learning and descriptive statistics. Prior works such as NoScope (Kang et al., 2017), TASTI (Kang et al., [n. d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML models over large datasets, typically in video analytics. Some optimizations proposed in these works, such as model cascades and predicate re-ordering, are also useful for optimizing LOTUS pipelines with language models. However, our bulk semantic processing setting requires new operators with significantly different semantics, such as sem_join and sem_groupby, which were not optimized in this prior work. These motivate our novel optimizations that can approximate these operators with statistical guarantees (Section 4).
Text2SQL A large body of prior work looks at Text2SQL (Yu et al., 2018; Yaghmazadeh et al., 2017; Zhang et al., 2024; Zelle and Mooney, 1996), using LMs to translate natural language queries to executable SQL code. These methods serve semantic queries that are expressible in traditioanl relational operators. By contrast, semantic operators provide new AI-based transformations, such as aggregations for summarizing large document sets, or rankings, joins and filters according to natural language expressions, which are not expressible in SQL. We note an interesting line of future work may build a natural language interface on top of semantic operators.
@@ 7. Conclusion
In this work, we proposed semantic operators to provide the first declarative and general-purpose interface to serve bulk-semantic processing. We implement these operators in the LOTUS system to seamlessly extend the relational model and allow programmers to easily compose powerful reasoning-based query pipelines over vast corpora of structured and unstructured data. Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of the semantic operator model as well as new opportunities for efficient algorithms and optimizations. For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized with accuracy guarantees to achieve higher performance than existing implementations.